Consider the following economy with Ta = 650 and YA = 9,800. a) If G = 3700, Derive the equation of the BDA as a function of t (tax rate). b) Find the tax rate at which the country will be running a budget balance (zero deficit). c) Find the values of the deficits for the following tax rates: 20%, 22%, 25%, 30%, 32%, 35% and 40%. d) Graph your findings in c using any graphing tool ( preferably excel) Please make sure you answer d and c. Need graph to show work
Answer #1

(a)

Total tax: T = Ta + t x YA

T = 650 + 9,800 x t

BDA (Budget deficit) = T - G

BDA = 650 + 9,800t - 3,700

BDA = 9,800t - 3,050

(b)

When budget is balanced, BDA = 0

9,800t - 3,050 = 0

9,800t = 3,050

t = 0.3112 = 31.12%

(c) [Note that when BDA > 0, there is budget deficit and when BDA < 0, there is budget surplus]

t BDA
0.2 -1090
0.22 -894
0.25 -600
0.3 -110
0.32 86
0.35 380
0.4 870

(d)

900 BDA 700 500 300 100 -100 02 0.3 0.25 0.35 0.4 -300 -500 700 006- Tax rate (t) -1100 BDA